### v1

@@ 2.3 Primordial black holes as dark matter
The gravitational wave detection by LIGO (Abbott et al., 2016) again raises interest in primordial black holes (PBH)s (Bird et al., 2016; Carr & Kühnel, 2020). The primordial black holes which were introduced by Zel’dovich and Novikov (Zel’dovich & Novikov, 1966) are a byproduct of the evolution of the Universe which can be produced in early universe (Escrivà, Kuhnel, & Tada, 2022). However, there are many sophisticated observational constraints on the abundance of primordial black holes in different ranges of their mass (Carr et al., 2017). In this work, we assume that a non-zero fraction of dark matter ($f_{\rm{PBH}}$) is PBHs.
In the realm of the standard model of cosmology and structure formation, the galaxies reside in DM halos (Mo, van den Bosch, & White, 2010). The distribution of DM halos nearly follows the Press-Schechter distribution consistent with the hierarchical structure formation (Press & Schechter, 1974). The N-body simulations and the observation of DM tracers on small scales show that the distribution of DM in halos is almost universal. To be more precise, the DM halo density profile nearly follows the one proposed by Navarro–Frenk–White (NFW) $\rho_{\text{\tiny{NFW}}}$ (Navarro et al., 1996):
@@
where $\rho_{s}$ and $r_{s}$ are the specific density and radius corresponding to each dark matter halo. The edge of each dark matter halo can be specified by its virial radius $r_{\rm{vir}}$, which can be related to the specific radius of the halo with concentration parameter $c$ as $r_{\rm{vir}}=cr_{s}$. The concentration parameter ranges from $404~{}\rm{to}~{}$ 40 for different DM halos. The total mass of dark matter halo $M_{\rm{halo}}$, also related to the specific radius and concentration as
@@
As the distribution of PBH in DM halo is almost uniform, the NFW distribution is sufficient to obtain the number density of PBH by the knowledge of $f_{\rm{PBH}}$ and the mass of PBH  [^6] In this work, we assume a monochromatic mass spectrum for PBH (Erfani et al., 2021; Kleban & Norton, 2023).. In the next section, we will show the correspondence of civilization type and the number of PBH used for harvesting the energy by ET-AI.
3 Search of Extraterrestrial-AI
In this section, we realize the idea of using the PBH in the DM halo of a galaxy as a source of energy for an AI-based civilization. Civilization type I uses the energy of its host star incident on the planet with an energy scale of $\sim 4\times 10^{16}$ W. Civilization type II uses the whole energy of the host star $\sim 4\times 10^{26}$ W and civilization type III uses whole energy of the galaxy $\sim 4\times 10^{37}$ W. For these scales, we assume a typical star-like sun in a typical galaxy such as Milky Way. Sagan introduced a civilization parameter due to the amount of energy consumption as (Sagan & Agel, 2000)
@@ 3 Search of Extraterrestrial-AI
where $P$ is the consumed power. In this conclusive section, we are in a position for searching extraterrestrial AI. We assume that they are living in the host planet-star system. However, they use energy storage from beyond their solar system for computation. Accordingly, we quantify the distance in which AI explores the space to collect the needed energy.
Now we define a specific length known as space exploration distance (sed) ${\cal{D}}_{\rm{sed}}$. We, human beings, as a civilization of $K\simeq 0.7$ (Sagan, 1973), install our facilities in Lagrangian point 2 as stable space construction, which exchanges the data. By stable construction, we mean structures that are in almost the same position with respect to the orbital position of the sun and Earth, like space telescopes in Lagrangian point 2. Besides them, we have already sent Voyager 1 &\&& 2 almost beyond the solar system. They are still functional and collecting data  [^7] For more information, see https://science.nasa.gov/mission/voyager . Inspired by this observation and an extrapolation, civilization Type I’s structure will reach scales of AU. So, type II will reach the gravitational boundaries of the solar system. We propose the approximate relation below between space exploration distance (sed) ${\cal{D}}_{\rm{sed}}$ and Sagan’s type of civilization:
@@
Our Civilization $1.7\times 10^{13}$ Watt $\sim 0.01$ AU
Type I $4\times 10^{16}$ Watt $\sim 1-2$ AU
Type II $4\times 10^{26}$ Watt $\sim 10^{5}$ AU
Type III $4\times 10^{37}$ Watt $\sim 4\times 10^{9}$ AU
Table 2: Civilization type in the first column is categorized versus the energy consumption in the second column, and Space Exploration distance (SED) ${\cal{D}}_{\rm{sed}}$ in the third column.
@@
To elaborate more, the equation 14 shows that for our current civilization stage the ${\cal{D}}_{\rm{sed}}(K\simeq 0.7)\simeq 0.01\rm{AU}$. As mentioned, we set the space exploration distance to the sun-earth Lagrangian point 2, where we set our astronomical instruments (space structures) there. Nowadays the European Space Agency’s (ESA) Gaia probe (Gaia Collaboration et al., 2016), the joint Russian-German high-energy astrophysics observatory Spektr-RG 888official webpage: https://www.srg.cosmos.ru/ which carries both extended Roentgen Survey with an Imaging Telescope Array (eROSITA) (Merloni et al., 2012) and Astronomical Roentgen Telescope X-ray Concentrator (ART-XC) (Pavlinsky et al., 2021), the joint National Aeronautics and Space Administration (NASA), ESA and Canadian Space Agency (CSA) James Webb Space Telescope (JWST) (Gardner et al., 2006) and the ESA’s Euclid mission (Laureijs et al., 2011).
For civilization type I, where one should harvest all the energy emitted from the planet, we set the astronomical unit scale. For civilization Type II, for harvesting all the emitted energy of the star, the scale is up to the gravitational scale of the solar system, which is the Oort cloud. For the final stage, we set the scale of the extension of the disk of the Milky Way as a typical galaxy. This approximation can relate the amount of power consumption of a civilization to its space exploration distance with a relation
@@
Now, we want to calculate the number of Dyson sphere-like structures due to the type of civilization. For this task, for a proof of concept, we assume a typical Milky-way galaxy with total mass of $M\simeq 1.5\times 10^{12}M_{\odot}$, the concentration of dark matter halo profile $c\simeq 10$ and the virial radius $r_{\rm{vir}}\simeq 200~{}\rm{kpc}$ (Grand et al., 2019). The fraction of primordial black holes (PBH) $f_{\rm{PBH}}$ and their mass are free parameters.
The number of PBH around the radius of space exploration distance is obtained
@@
where $r_{*}\simeq 8~{}{\rm{kpc}}$ is the position of the planet in Milky Way. Here, we assume that the life-harbouring planet whose inhabitants are AI resides far away from the centre of the galaxy. However, as we know, the distribution in dark matter halo is spherical with a good approximation. The number of PBHs are
@@
where ${\cal{F}}(r_{*},{\cal{D}}_{\rm{sed}};c)$ is a function of the concentration of DM halo and distance of the planet from the centre of the galaxy and space extrapolation distance as below
@@
In figure 1, we plot a contour representation of the chance that an advanced civilization will provide its whole energy from our proposal. It means that for a certain type of Sagan parameter $K$, the energy is provided only by the Dyson-sphere-like structures around PBH. The $x$-axis is Sagan’s parameter of the Kardashev civilization type, which determines the energy needed for a civilization. The $y$-axis is the fraction of PBH for a mass of $M_{\rm{\tiny{PBH}}}=1M_{\odot}$. The colour bar of the contours shows the quantity $K_{\rm{achieved}}/K-1$, in which $K_{\rm{achieved}}$ is the amount of civilization parameter achieved from our proposed method using only PBHs. The bluish regions (with negative values) are not suitable for our proposal. Note that the contour plot is obtained by following steps: First, by knowing the Sagan parameter $K$ and equation 14, we can find the space exploration length. In the next step, by using equations 17 and 18, we can find the number of PBH from the host planet of ET-AI up to the scale of ${\cal{D}}_{\rm{sed}}$. Now, in the case, $K\leq K_{\rm{achieved}}$ means all energy can be obtained from PBH. Figure 1 shows that for a specific $K$ value, by increasing the fraction of PBH, $f_{\rm{PBH}}$, we can achieve more energy (corresponding to $K_{\rm{achieved}}$). For a specific fixed $f_{\rm{PBH}}$, the chance of providing all the energy from PBH decreases by a larger Sagan scale. This is a consequence of the fact that the slope of needed energy is higher than the energy which ET-AI can obtain from PBH at larger distances.
In figure 2, for a $f_{\rm{\tiny{PBH}}}=0.01$, we plot the contours of $K_{\rm{achieved}}/K-1$ for primordial black with mass in the range of $0.01-100M_{\odot}$ as the source of civilization with Sagan $K$ parameter. Again, we should note that $K_{\rm{achieved}}$ is obtained by the method we describe for figure 1. The plot indicates that for a specific $K$ (space exploration length), the amount of achievable energy from our method is almost independent of the mass of the PBH. We anticipate this as the amount of energy from a PBH accretion disk is proportional to its mass. On the other hand, the abundance of smaller PBH compromises the amount of energy each can deliver to a Dyson sphere-like structure. The two plots show that the main parameter is the fraction of PBH.
Figure 1: The $x$-axis is the Sagan’s parameter of the Kardashev civilization type. The $y$-axis is the fraction of PBH for a mass of $M_{\rm{\tiny{PBH}}}=1M_{\odot}$. And the colour-bar of the contours shows the quantity $K_{\rm{achieved}}/K-1$, which $K_{\rm{achieved}}$ is the amount of civilization parameter achieved from PBHs.
@@
The figures show that there is a reasonable chance that in a proper mass range of PBH and the non-zero fraction of PBH as dark matter, AI civilization uses PBH as the source of energy. Note that we choose a conservative value for $\eta_{\rm{disk}}$.
To quantify this we provide a numerical expression for the minimum fraction of PBH ($f_{\rm{PBH|min}}$) to satisfy the condition $K_{\rm{achieved}}/K-1>0$, (the boundary between the positive and negative regions of contour plot in figure 1 ), as a function of $K$:
@@
In terms of space exploration length, this relation will be:
@@
Note that we normalize the space exploration length to $1\rm{pc}$.
@@
Now, the question is about the observational consequences of SET-AI. The previous ideas are based on the technique for searching Dyson spheres around stars in the Milky Way. The idea is to look at the Infrared excess radiation of stars. A specific example of this kind of study appeared in works of Suazo et al. (2022); Contardo & Hogg (2024). The Infrared excess can be an effect of circumstellar dust, protoplanetary disks, or debris disks, and it has a degeneracy with a Dyson-like structure.
@@
For our proposal, we suggest that ET-AI uses PBHs as the energy source. Observing an excess in the submillimeter and Infrared radiation in the interstellar medium could indicate the Dyson sphere. For an advanced civilization which brings the Dyson sphere to a temperature $10~{}\rm{K}$ with limited efficiency, the waste energy will be in a black body radiation in $\sim 0.3~{}\rm{mm}$ from a structure with an extension of an order of $\sim 2700~{}\rm{AU}$ around a PBH accretion disk. We obtain this scale by equation 9, for a specific case where we assume a PBH mass of $1M_{\odot}$ and efficiency of $\eta_{\rm{disk}}=10^{-4}$.
@@
This sub-millimetre excess can be observed by the Atacama Large Millimeter/submillimeter Array (ALMA) by its ”band 10” detector working in $\sim 950~{}\rm{GHz}$ designed by the National Astronomical Observatory of Japan (Uzawa et al., 2013). This means that this excess can resemble a multi-blackbody spectrum akin to an accretion disk of PBH (see figure 3 in Hsiao et al. (2021)). Their results are based on multicolour blackbody radiation calculation of a disk with a Dyson sphere (see Mitsuda et al. (1984); Makishima et al. (1986); Merloni, Fabian, & Ross (2000)). However, in our scenario, the two peaks of the black body are separated further in wavelengths. For the accretion disk, one should observe the spectrum in X-ray and UV, while for the Dyson sphere, one should go to sub-millimetre scales. On the other hand, the spatial resolution of ALMA in band 10 is $\sim 0.5^{\prime\prime}$. This means a resolution of $\sim 2.42\times 10^{-6}\rm{rad}$. Accordingly, for a $\sim 2700\rm{AU}$ extended space structure, for direct imaging, we can explore our galaxy up to the scales $\sim 5.4~{}\rm{kpc}$. However, this technique needs a concrete study closely tied to ALMA’s sensitivity, depth, and required exposure time for the detection, which is out of the scope of the current work. To distinguish these images from other celestial objects, one has to refer to the spectroscopy of the Dyson sphere candidate as well. For solar mass PBH with $\eta_{\rm{disk}}\sim 10^{-4}$ and efficiency of $30\%$ for the Dyson sphere, we will have a solar Luminosity $\sim L_{\odot}$ in submillimeter in a black body radiation in angular resolution of $(\frac{2700\rm{AU}}{x_{\rm{PBH}}})\rm{rad}$, where $x_{\rm{PBH}}$ is the distance from PBH in astronomical units. This excess in blackbody shape must be compared with the same wavelength radiation in the interstellar medium for detection. Also, we anticipate an excess in higher frequencies due to the accretion disk at that region. In other words, the Dyson sphere-like structure radiation is a signal we want to extract from the interstellar radiation in the same wavelength, working as a noise for this specific detection.
@@
Another observational consequence of our work is the quest for the SET-AI in the extragalactic scales. This means that the galaxies will be in a state of excess radiation in infrared and submillimeter for the scales proportional to space exploration distance. Accordingly, one target of the future SET-AI will be galaxies with considerable Infrared radiation. For a specific example to consider: Assume an advanced civilization in Sagan scale $K\simeq 2.2$, with the assumption of a fraction of the $f_{\rm{PBH}}=1.0$, in which the idea of using PBH as a source of energy is achievable. The space exploration length in this configuration is in order of $\sim 4.8{\rm{pc}}$. It means that with ALMA band 10 with a given resolution of $\sim 0.5^{\prime\prime}$, we can probe the nearby galaxies up to $\sim 2~{}{\rm{Mpc}}$ for both the excess in $\sim 0.3~{}\rm{mm}$ spectrum and direct imaging. The wavelength mentioned again is suitable for Dyson sphere-like structures in $\sim 10\rm{K}$. For galaxies in cosmological distances, one can refer to an interesting way of quantifying the detection of the Infrared excess. It is the use of the excess Infrared radiation in a plot versus the UV-continuum slope. For more details see Bouwens et al. (2016). In distant galaxies, the angular resolution is not enough for AI-designed structure detection. However, one can find a sign of ET-AI for further investigation by excess radiation in higher redshifts.

### v2

@@ 2.3 Primordial black holes as dark matter
Gravitational wave detections by LIGO (Abbott et al., 2016) again raise interest in PBHs (Bird et al., 2016; Carr & Kühnel, 2020). PBHs which were introduced by Zel’dovich and Novikov (Zel’dovich & Novikov, 1966) are a by-product of the evolution of the Universe, which can be produced in the early Universe (Escrivà, Kuhnel, & Tada, 2022). However, there are many sophisticated observational constraints on the abundance of PBHs in different ranges of their mass (Carr et al., 2017). In this work, we assume that a nonzero fraction of DM ($f_{\rm{PBH}}$) is PBHs.
In the realm of the standard model of cosmology and structure formation, galaxies reside in DM halos (Mo, van den Bosch, & White, 2010). The distribution of DM halos nearly follows a Press-Schechter distribution consistent with hierarchical structure formation (Press & Schechter, 1974). $N$-body simulations and observations of DM tracers on small scales show that the distribution of DM in halos is almost Universal. To be more precise, the DM halo density profile nearly follows the one proposed by Navarro–Frenk–White (NFW), $\rho_{\text{\tiny{NFW}}}$ (Navarro et al., 1996):
@@
where $\rho_{s}$ and $r_{s}$ are, respectively, the specific density and radius corresponding to each DM halo. The edge of each DM halo can be specified by its virial radius $r_{\rm{vir}}$, which can be related to the specific radius of the halo with concentration parameter $c$ as $r_{\rm{vir}}=cr_{s}$. The concentration parameter ranges from $4-40$ for different DM halos. The total mass of the DM halo $M_{\rm{halo}}$, also related to the specific radius and concentration as
@@
As the distribution of PBHs in a DM halo is almost uniform, the NFW distribution is sufficient to obtain the number density of PBHs via the knowledge of $f_{\rm{PBH}}$ and the total mass of PBHs  [^6] In this work, we assume a monochromatic mass spectrum for PBHs (Erfani et al., 2021; Kleban & Norton, 2023).. In the next section, we will show the correspondence of civilization type and the number of PBH used for harvesting the energy by ET-AI.
3 Search for Extraterrestrial Artificial Intelligence
In this section, we realize the idea of using PBHs in the DM halo of a galaxy as a source of energy for an AI-based civilization. A type I civilization uses the energy of its host star that is incident on the planet with an energy scale of $\sim 4\times 10^{16}$ W. A type II civilization uses the whole energy of the host star, $\sim 4\times 10^{26}$ W and a type III uses whole energy of the galaxy, $\sim 4\times 10^{37}$ W. For these scales, we assume a typical star like the Sun in a typical galaxy such as the Milky Way. Sagan introduced a civilization parameter due to the amount of energy consumption as (Sagan & Agel, 2000)
@@ 3 Search for Extraterrestrial Artificial Intelligence
where $P$ is the consumed power. In this conclusive section, we are in the position to search for ET-AI. We assume that they are living in a host planet-star system. However, they use energy storage from beyond their solar system for computation. Accordingly, we quantify the distance in which AI explores the space to collect the needed energy.
Now we define a specific length known as SED ${\cal{D}}_{\rm{sed}}$. We, human beings, as a civilization of $K\simeq 0.7$ (Sagan, 1973), install our facilities in the Lagrangian 2 point as a stable space construction, which exchanges the data. By stable construction, we mean structures that are in almost the same position with respect to the orbital position of the sun and Earth, like space telescopes at the Lagrangian 2 point. Besides them, we have already sent Voyager 1 and 2 almost beyond the solar system. They are still functional and collecting data  [^7] For more information, see https://science.nasa.gov/mission/voyager . Inspired by this observation and an extrapolation, a type I civilization’s structure will reach astronomical unit scales. So, type II civilizations will reach the gravitational boundaries of their solar system. We propose the approximate relation below between SED ${\cal{D}}_{\rm{sed}}$ and Sagan’s type of civilization
@@
Our Civilization $1.7\times 10^{13}$ W $\sim 0.01$ AU
Type I $4\times 10^{16}$ W $\sim 1-2$ AU
Type II $4\times 10^{26}$ W $\sim 10^{5}$ AU
Type III $4\times 10^{37}$ W $\sim 4\times 10^{9}$ AU
Table 2: Civilization Type in the first column is categorized vs. the energy consumption in the second column and Space Exploration distance ${\cal{D}}_{\rm{sed}}$ in the third column.
@@
To elaborate more, the equation 14 shows that for our current civilization stage the ${\cal{D}}_{\rm{sed}}(K\simeq 0.7)\simeq 0.01\rm{AU}$. As mentioned, we set the SED to the Sun-Earth Lagrangian 2 point, where we set our astronomical instruments (space structures). Nowadays the European Space Agency’s (ESA) Gaia probe (Gaia Collaboration et al., 2016); the joint Russian–German high-energy astrophysics observatory Spektr–RG, 888official webpage: https://www.srg.cosmos.ru/ which carries both the extended Roentgen Survey with the Imaging Telescope Array (eROSITA) (Merloni et al., 2012) and Astronomical Roentgen Telescope X-ray Concentrator (ART-XC) (Pavlinsky et al., 2021); the joint National Aeronautics and Space Administration, ESA and Canadian Space Agency (CSA) James Webb Space Telescope (JWST) (Gardner et al., 2006) and the ESA’s Euclid mission (Laureijs et al., 2011), are all at the Lagrangian 2 point.
For a type I civilization, where all of the energy emitted from the planet is harvested, we use the astronomical unit scale. For a type II civilization, where all of the emitted energy of the star is harvested, the scale is up to the gravitational scale of the solar system, which is the Oort cloud. For the final stage, we set the scale of the extension of the disk of the Milky Way as a typical galaxy (see Table 2). This approximation can relate the amount of power consumed by a civilization to its SED with the relation
@@
Now, we want to calculate the number of Dyson sphere–like structures as a function of the type of civilization. For this task, for a proof of concept, we assume a typical Milky Way galaxy with a total mass of $M\simeq 1.5\times 10^{12}M_{\odot}$, the concentration of the dark matter halo profile $c\simeq 10$ and virial radius $r_{\rm{vir}}\simeq 200~{}\rm{kpc}$ (Grand et al., 2019). The fraction of PBHs $f_{\rm{PBH}}$ and their mass are free parameters.
The number of PBH around the radius of the SED is obtained as
@@
where $r_{*}\simeq 8~{}{\rm{kpc}}$ is the position of the planet in Milky Way. Here, we assume that the life-harboring planet whose inhabitants are AI resides far away from the center of the galaxy. However, as we know, the distribution in DM halo is spherical with a good approximation. The number of PBHs is
@@
where ${\cal{F}}(r_{*},{\cal{D}}_{\rm{sed}};c)$ is a function of the concentration of DM halo and distance of the planet from the center of the galaxy and SED as below
@@
In figure 1, we plot a contour representation of the chance that an advanced civilization will provide its whole energy from our proposal. It means that for a certain type of Sagan’s parameter $K$, the energy is provided only by the Dyson sphere–like structures around PBHs. The $x$-axis is Sagan’s parameter of the Kardashev civilization type, which determines the energy needed for a civilization. The $y$-axis is the fraction of PBH for a mass of $M_{\rm{\tiny{PBH}}}=1M_{\odot}$. The color bar of the contours shows the quantity $K_{\rm{achieved}}/K-1$, in which $K_{\rm{achieved}}$ is the amount of civilization parameter achieved from our proposed method using only PBHs. The bluish regions (with negative values) are not suitable for our proposal. Note that the contour plot is obtained by following steps: First, by knowing the Sagan’s parameter $K$ and equation 14, we can find the space exploration length. In the next step, by using equations 17 and 18, we can find the number of PBH from the host planet of ET-AI up to the scale of ${\cal{D}}_{\rm{sed}}$. Now, the case, $K\leq K_{\rm{achieved}}$ means all energy can be obtained from PBHs. Figure 1 shows that for a specific $K$ value, by increasing the fraction of PBH, $f_{\rm{PBH}}$, more energy can be provided (corresponding to $K_{\rm{achieved}}$). For a specific fixed $f_{\rm{PBH}}$, the chance of providing all of the energy from PBHs decreases at larger Sagan scales. This is a consequence of the fact that the amount of needed energy is higher than the energy which ET-AI can obtain from PBHs at larger distances.
In figure 2, for a $f_{\rm{\tiny{PBH}}}=0.01$, we plot the contours of $K_{\rm{achieved}}/K-1$ for PBHs with mass in the range of $0.01-100M_{\odot}$ as a function of Sagan’s $K$ parameter. Again, we should note that $K_{\rm{achieved}}$ is obtained by the method we describe for figure 1. The plot indicates that for a specific $K$ (space exploration length), the amount of achievable energy from our method is almost independent of the mass of the PBH. We anticipate this as the amount of energy from a PBH accretion disk is proportional to its mass. On the other hand, the abundance of smaller PBHs compromises the amount of energy each can deliver to a Dyson sphere–like structure. The two plots show that the main parameter is the fraction of PBHs.
Figure 1: The $x$-axis is Sagan’s parameter of the Kardashev civilization type. The $y$-axis is the fraction of PBH for a mass of $M_{\rm{\tiny{PBH}}}=1M_{\odot}$. The color bar of the contours shows the quantity $K_{\rm{achieved}}/K-1$, in which $K_{\rm{achieved}}$ is the amount of civilization parameter achieved from PBHs.
@@
The figures show that there is a reasonable chance that in a proper mass range of PBHs and a nonzero fraction of PBHs serving as dark matter, an AI civilization uses PBHs as a source of energy. Note that we choose a conservative value for $\eta_{\rm{disk}}$.
To quantify this we provide a numerical expression for the minimum fraction of PBH ($f_{\rm{PBH|min}}$) to satisfy the condition $K_{\rm{achieved}}/K-1>0$, (the boundary between the positive and negative regions of contour plot in figure 1 ), as a function of $K$.
@@
In terms of space exploration length, this relation will be
@@
Note that we normalize the space exploration length to $1\rm{pc}$. Now, the question is about the observational consequences of the SET-AI. The previous ideas are based on the technique for searching for Dyson spheres around stars in the Milky Way. The idea is to look at the infrared excess radiation of stars. A specific example of this kind of study appeared in works of Suazo et al. (2022); Contardo & Hogg (2024). The Infrared excess can be an effect of circumstellar dust, protoplanetary disks, or debris disks, and it has a degeneracy with a Dyson sphere–like structure.
@@
For our proposal, we suggest that ET-AI uses PBHs as an energy source. Observing an excess in the submillimeter and infrared radiation in the interstellar medium could indicate a Dyson sphere. For an advanced civilization which brings a Dyson sphere to a temperature of $10~{}\rm{K}$ with limited efficiency, the waste energy will be in a blackbody radiation at $\sim 0.3~{}\rm{mm}$ from a structure with an extension of the order of $\sim 2700~{}\rm{AU}$ around a PBH accretion disk. We obtain this scale by equation 9, for a specific case where we assume a PBH mass of $1M_{\odot}$ and an efficiency of $\eta_{\rm{disk}}=10^{-4}$.
@@
This submillimetre excess can be observed by the Atacama Large Millimeter/submillimeter Array (ALMA) by its Band 10 detector working at $\sim 950~{}\rm{GHz}$, as designed by the National Astronomical Observatory of Japan (Uzawa et al., 2013). This means that this excess can resemble a multiblackbody spectrum akin to the accretion disk of a PBH (see figure 3 in Hsiao et al. (2021)). Their results are based on a multicolour blackbody radiation calculation of a disk with a Dyson sphere (see Mitsuda et al. (1984); Makishima et al. (1986); Merloni, Fabian, & Ross (2000)). However, in our scenario, the two peaks of the blackbody are separated further in wavelength. For the accretion disk, one should observe the spectrum in X-ray and UV, while for the Dyson sphere, one should go to submillimetre scales. On the other hand, the spatial resolution of ALMA in Band 10 is $\sim 0.5^{\prime\prime}$. This means a resolution of $\sim 2.42\times 10^{-6}\rm{rad}$. Accordingly, for a $\sim 2700\rm{AU}$ extended space structure, for direct imaging, we can explore our Galaxy up to scales $\sim 5.4~{}\rm{kpc}$. However, this technique needs a concrete study closely tied to ALMA’s sensitivity, depth, and required exposure time for a detection, which is out of the scope of the current work. To distinguish these images from other celestial objects, one has to get spectroscopy of the Dyson sphere candidate as well. For a solar–mass PBH with $\eta_{\rm{disk}}\sim 10^{-4}$ and efficiency of $30\%$ for the Dyson sphere, we will have solar Luminosity $\sim 1L_{\odot}$ at submillimeter in a blackbody radiation with angular resolution of $(\frac{2700\rm{AU}}{x_{\rm{PBH}}})\rm{rad}$, where $x_{\rm{PBH}}$ is the distance from the PBH in astronomical units. This excess in blackbody shape must be compared with the same wavelength radiation in the interstellar medium for detection. Also, we anticipate an excess at higher frequencies due to the accretion disk at that region. In other words, the Dyson sphere–like structure radiation is a signal we want to extract from measurement of interstellar radiation while accounting for noise at the same wavelength.
@@
Another observational consequence of our work is the quest for the SET-AI at the extragalactic scales. This means that the galaxies will be in a state of excess radiation in the infrared and submillimeter for the scales proportional to SED. Accordingly, one target of the future SET-AI will be galaxies with considerable infrared radiation. For a specific example let us consider the following. Assume an advanced civilization with Sagan’s scale $K\simeq 2.2$, with the assumption of a fraction of the $f_{\rm{PBH}}=1.0$, in which the idea of using PBHs as a source of energy is achievable. The space exploration length in this configuration is in order of $\sim 4.8{\rm{pc}}$. It means that with ALMA Band 10 with a given resolution of $\sim 0.5^{\prime\prime}$, we can probe the nearby galaxies up to $\sim 2~{}{\rm{Mpc}}$ for excess at $\sim 0.3~{}\rm{mm}$ is spectra and direct imaging. The wavelength mentioned again is suitable for Dyson sphere–like structures with $\sim 10\rm{K}$. For galaxies at cosmological distances, one can refer to an interesting way of quantifying the detection of the infrared excess: looking for excess infrared radiation in a plot versus the UV continuum slope. For more details see Bouwens et al. (2016). In distant galaxies, the angular resolution is not enough for AI-designed structure detection. However, one can find a sign of ET-AI for further investigation by excess radiation in higher redshifts.